OpenStudy (jacksonjrb):

Algebraically determine the domain and range: y=x^2-8x+7

OpenStudy (jacksonjrb):

i already know the domain just need help with the range

jimthompson5910 (jim_thompson5910):

have you learned about completing the square?

OpenStudy (jacksonjrb):

Yes

OpenStudy (anonymous):

wait you got x you said right?

OpenStudy (jacksonjrb):

\[x=y^2-8y+7\]

OpenStudy (jacksonjrb):

And yes

jimthompson5910 (jim_thompson5910):

complete the square to get the equation into vertex form y = a(x-h)^2 + k the range will be \(\Large y \ge k\) (replace k with a numeric value though) because 'a' is positive which means the parabola opens upward

OpenStudy (jacksonjrb):

I was taught to substitute the x for the y and vice vera

OpenStudy (anonymous):

@jim_thompson5910 can't he use substitution?

jimthompson5910 (jim_thompson5910):

@JacksonJRB you're thinking of the inverse

OpenStudy (jacksonjrb):

Ah

jimthompson5910 (jim_thompson5910):

first compute h = -b/(2a) in this case, a = 1, b = -8

OpenStudy (freckles):

there might be an easier way x^2-8x+7 is factorabole y=x^2-8x+7 you can find the average of the 0's so the 0's are a and b the average of a and b is (a+b)/2 so you can find the range by doing: \[[f(\frac{a+b}{2}),\infty) \text{ since } a=1>0\]

jimthompson5910 (jim_thompson5910):

once you know the value of h, plug it into y=x^2-8x+7 to find k

OpenStudy (solomonzelman):

y-value of the vertex (the k of the vertex) is the minimum range. (and there is no maximum limit for range) no range limits, since it is a polynomial

jimthompson5910 (jim_thompson5910):

`can't he use substitution?` I'm not sure what you mean @GTA_Hunter35

OpenStudy (anonymous):

he said he has x

jimthompson5910 (jim_thompson5910):

He said he had the domain. Not just a single value of x.

OpenStudy (anonymous):

nvm it won't work

OpenStudy (jacksonjrb):

^^

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

its your call @jim_thompson5910

OpenStudy (jacksonjrb):

I'm still very confused...

jimthompson5910 (jim_thompson5910):

were you able to compute h = -b/(2a) ?

OpenStudy (freckles):

do you want to complete the square or find the zeros or use the vertex formula?

OpenStudy (jacksonjrb):

Yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what value did you get for h

OpenStudy (anonymous):

wait wats the vertex formula?

OpenStudy (jacksonjrb):

And either way @freckles

OpenStudy (jacksonjrb):

h=4

OpenStudy (freckles):

well if you find the average of the zeros you will find the x-coordinate of the vertex

jimthompson5910 (jim_thompson5910):

plug x = 4 into y=x^2-8x+7 and you get y = ??

OpenStudy (freckles):

nevermind you guys got the x-coordinate of the vertex

OpenStudy (jacksonjrb):

-9

jimthompson5910 (jim_thompson5910):

so k = -9 making the range \(\Large y \ge -9\) we have a parabola that has the lowest point at (4,-9). All other points will have larger y values.

OpenStudy (jacksonjrb):

Ah. That makes sense now. Thank you so much!

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (freckles):

\[f(x)=x^2-8x+7 \\ f(x)=(x-7)(x-1) \\ f(x)=0 \text{ when } x=1 \text{ or } x=7 \\ \text{ the average of the zeros is } \frac{7+1}{2}=\frac{8}{2}=4 \\ \text{ so since } a=1>0 \text{ then the range is } [f(4),\infty) \\ f(4)=4^2-8(4)+7=-9 \\ \text{ so the range is } [-9,\infty)\] just wanted to type what I was going for

OpenStudy (anonymous):

well ur right too but u used the function way

OpenStudy (freckles):

used the function way?

OpenStudy (anonymous):

u used f of x

OpenStudy (anonymous):

dat was the only difference

OpenStudy (freckles):

k